OpenStudy (anonymous):

Which calculation could be used to determine the area of this square? A. A B. A C. A D. A

OpenStudy (anonymous):

A= A=7+7

OpenStudy (anonymous):

B. A=7+7+7+7

OpenStudy (anonymous):

C. A=7X4

OpenStudy (anonymous):

D. A=7X7

OpenStudy (anonymous):

Help @ace-n-it

OpenStudy (deathlynine):

A = a^2

OpenStudy (anonymous):

um C.

OpenStudy (deathlynine):

Each side of the square is 7. Try again?

OpenStudy (anonymous):

a.

OpenStudy (deathlynine):

Area of square is being multiplying so A and B are crossed out.

OpenStudy (anonymous):

D.

OpenStudy (deathlynine):

Correct! :)

OpenStudy (anonymous):

thank you i passed